A pulley is set up on an incline surface. The incline surface has a static coefficient u s and kinetic u k. m2 moves downhill when the pulley is free.

The pulley has a moment of inertia I = -mr 2 and a mass M.

T M R m 2 m1

a. Find angular acceleration of the pulley.

b. Find the tension resulting from the two masses in the string.

c. Find the reaction forces at the pulley
Answer #1

Solution : . MIR T2 Mg, uk a fe Is 12mp² B Magsing mig magloss Fon mi To mig=mia - (0) For magsinß- fr = mza - Tz -(10) becaufrom equation () Tormiga mig (ma Sinß- uk M2 Co BB- mi) (+7+2) Te migl m2. Sins - Ukmecos B-mi mitm24 m2 and using To-T1=Ma T